# Structure-preserving and thermodynamically consistent finite element discretization for visco-resistive MHD with thermoelectric effect

### v1

@@ 1 Introduction
Plasma flows are at the core of numerous applications, including but not limited to: astrophysics, Magnetic Confinement Fusion [5, 25] and Inertial Confinement Fusion [1, 6]. Such flows display complex behaviors because of thermodynamics, interaction with the electromagnetic field, potentially highly compressible effects, numerous and stiff irreversible processes. In most applications, characteristic velocities are negligible compared to the speed of light. In this case, the flow can be reasonably approximated by Magnetohydrodynamics (MHD). The equations under interest are as follows
@@
with $n$ the outward normal vector to the boundary, together with a thermal boundary condition of either Dirichlet or Neumann type. Extension of these equations to include thermoelectric effect will be also considered. When dissipation (i.e. viscosity, resistivity and heat exchange) is removed, the so-called ideal MHD is recovered. It is well-known that ideal MHD equations can be derived from Hamilton’s principle [24]. Because dissipation is by nature irreversible and leads to an entropy production, it falls in the realm of non-equilibrium thermodynamics. An extension of Hamilton’s principle for non-equilibrium thermodynamics was developed in [13, 14]. It allows to derive equations automatically satisfying the two first laws of thermodynamics; namely conservation of energy and non-destruction of entropy. The above equations (1), including all irreversible processes, can be obtained from it, as will be shown shortly. This structure helps navigate through the otherwise complicated and intricate phenomena. More importantly, it will here serve as a guideline for deriving a numerical method that preserves key properties of the flow at the discrete level.
@@ 2.1 Ideal MHD
In the reversible case, as a consequence of mass and entropy conservation, and of the frozen-in property of magnetic induction, one has $\varrho(t)=\varrho_{0}$, $S(t)=S_{0}$ and $\mathcal{B}(0)=\mathcal{B}_{0}$ in the Lagrangian description. In this case, the MHD motion follows from Hamilton’s principle
@@
The 00-form $\gamma$ and $1$-form $Z$ will be used later in the irreversible case. The reduced Hamilton’s principle (4) then becomes
@@
$\partial_{t}\frac{\delta\ell}{\delta u}+\pounds_{u}\frac{\delta\ell}{\delta u}$ $=\rho\nabla\frac{\delta\ell}{\delta\rho}+s\nabla\frac{\delta\ell}{\delta s}+B\times\operatorname{curl}\nabla\frac{\delta\ell}{\delta B},$ (13a)
@@ 2.2 Visco-resistive MHD with heat conduction and thermoelectric effect
$\partial_{t}\frac{\delta\ell}{\delta u}+\pounds_{u}\frac{\delta\ell}{\delta u}=\rho\nabla\frac{\delta\ell}{\delta\rho}+s\nabla\frac{\delta\ell}{\delta s}+\operatorname{div}\sigma+B\times\operatorname{curl}\frac{\delta\ell}{\delta B}$ (22a)
$\overline{\mathfrak{D}}_{t}B=\operatorname{curl}j_{B}$ (22b)
$\overline{D}_{t}\rho=0$ (22c)
$\displaystyle\frac{\delta\ell}{\delta s}(\overline{D}_{t}s+\operatorname{div}j_{s})=-\sigma:\nabla u-j_{s}\cdot\nabla\frac{\delta\ell}{\delta s}-j_{B}\cdot\operatorname{curl}\frac{\delta\ell}{\delta B}$ (22d)
@@
$j_{s}\cdot n=0,\quad j_{B}\times n=0$ (23)
@@
The variational formulation given above yields thermally insulated boundaries, as we see from the first equation in (23), consistently with the variational setting of adiabatically closed systems, [13]. As shown in [11], a modification of the constraints (20a) and (20b), consistent with the variational principle for open systems, [15], allows the treatment of the Dirichlet boundary condition $T|_{\partial\Omega}=T_{0}$ or the nonhomogeneous Neumann boundary condition $j_{q}\cdot n=q_{0}$, with $j_{q}=Tj_{s}$ the heat flux, for given $T_{0},q_{0}:\partial\Omega\rightarrow\mathbb{R}$. To obtain these boundary conditions, the constraint (20a) is modified with the addition of a boundary term
$\int_{\Omega}w\frac{\delta\ell}{\delta s}\overline{D}_{t}\varsigma\,{\rm d}x$ $=\int_{\Omega}w\big{(}-\sigma:\nabla u+j_{s}\cdot\nabla D_{t}\gamma+j_{B}\cdot\operatorname{curl}(\mathfrak{D}_{t}Z)\big{)}\,{\rm d}x-\int_{\partial\Omega}\!\!\!w(j_{s}\cdot n)\left(D_{t}\gamma-T_{0}\right)\,{\rm d}a,\;\forall\;w$ (27a)
$\int_{\Omega}w\frac{\delta\ell}{\delta s}\overline{D}_{t}\varsigma\,{\rm d}x$ $=\int_{\Omega}w\big{(}-\sigma:\nabla u+j_{s}\cdot\nabla D_{t}\gamma+j_{B}\cdot\operatorname{curl}(\mathfrak{D}_{t}Z)\big{)}\,{\rm d}x-\int_{\partial\Omega}w\big{(}(j_{s}\cdot n)D_{t}\gamma-q_{0}\big{)}\,{\rm d}a,\;\forall\;w$ (27b)
for the Dirichlet boundary conditions, and for the nonhomogeneous Neumann boundary condition. For both cases, the variational constraint (20b) is modified as
@@
The variational principle (19)–(21) with (20a) replaced by (27a) or (27b), and with (20b) replaced by (28), yield the same system (22)–(24), but with the boundary condition $j_{s}\cdot n=0$ replaced by $T|_{\partial\Omega}=T_{0}$ or $(j_{s}\cdot n)T=q_{0}$. In these cases, the energy balance in (25) is modified as
@@ 2.3 Physical closures
with $\varepsilon(\rho,s)$ the internal energy density and $\mu_{0}>0$ the magnetic permeability. One then gets $\frac{\delta\ell}{\delta u}=\rho(u+R)$, $\frac{\delta\ell}{\delta\rho}=\frac{1}{2}|u|^{2}+R\cdot u-\frac{\partial\epsilon}{\partial\rho}-\phi$, $-\frac{\delta\ell}{\delta s}=\frac{\partial\epsilon}{\partial s}=T>0$ the temperature, and $-\frac{\delta\ell}{\delta B}=\frac{1}{\mu_{0}}B=H$. By defining the pressure $p=\frac{\partial\epsilon}{\partial\rho}\rho+\frac{\partial\epsilon}{\partial s}s-\epsilon$, (22a) and (22d) eventually become
$\rho(\partial_{t}u+u\cdot\nabla u+2\omega\times u)$ $=-\rho\nabla\phi-\nabla p+\operatorname{curl}H\times B+\operatorname{div}\sigma$ (31a)
@@
$j_{s}$ $=j_{s}(T)=-\frac{1}{T}\kappa\nabla T,$ with $\text{with $\kappa\geq 0$},$ (32b)
$j_{B}$ $=j_{B}(H)=-\nu\operatorname{curl}H,$ with $\text{with $\nu\geq 0$},$ (32c)
@@
$j_{s}$ $=j_{s}(T,H)=-\frac{1}{T}\kappa\nabla T+\alpha\operatorname{curl}H,$ with $\text{with $\kappa\geq 0$},$ (34a)
$j_{B}$ $=j_{B}(T,H)=-\alpha\nabla T-\nu\operatorname{curl}H,$ with $\text{with $\nu\geq 0$},$ (34b)
for some coefficient $\alpha$. Contrary to the previous coefficients, $\alpha$ is not subject to any sign restriction, as the thermoelectric effect only contributes to the entropy balance through a flux, the so called Thomson effect. Note that if $\alpha$ is constant or a sole function of the temperature, it does not contribute to the induction equation.
@@ 2.4 Associated weak formulation
$e(w,f)=\left\{\begin{array}[]{ll}\vspace{0.2cm}\displaystyle 0&\text{for homogeneous Neumann}\\ \vspace{-0.4cm}\displaystyle\int_{\partial\Omega}wj_{s}(f)\cdot nT_{0}\,{\rm d}a&\text{for Dirichlet}\\ \\ \displaystyle\int_{\partial\Omega}wq_{0}(f)\,{\rm d}a&\text{for nonhomogeneousNeumann}.\end{array}\right.$ (37)
@@
With these definitions, the constraints (20a)–(20b), (27a)–(28), and (27b)–(28) can be written in a unified way for all boundary conditions as
@@
[…]
@@
By using these notations when carrying out the variational formulations (19)–(21) or (19)-(27a)-(28)-(21) or (19)-(27b)-(28)-(21), we get the following weak form of the equations (22) and associated boundary conditions $j_{s}\cdot n=0$ or $T|_{\partial\Omega}=T_{0}$ or $j_{q}\cdot n=q_{0}$ as follows
@@ 3.1 Discrete setting
$H^{1}_{0}(\Omega)$ =$=\{f\in L^{2}(\Omega)\,|\,\nabla f\in L^{2}(\Omega)^{3},\,f=0\text{ on }\partial\Omega\},$ (41a)
$H_{0}(\operatorname{div},\Omega)$ =$=\{u\in L^{2}(\Omega)^{3}\,|\operatorname{div}u\in L^{2}(\Omega),\,u\cdot n=0\text{ on }\partial\Omega\},$ (41b)
$H_{0}(\operatorname{curl},\Omega)$ =$=\{u\in L^{2}(\Omega)^{3}\,|\,\operatorname{curl}u\in L^{2}(\Omega),\,u\times n=0\text{ on }\partial\Omega\}.$ (41c)
@@
$U^{\text{grad}}_{h}=CG_{r+1}(\mathcal{T}_{h})^{3}$ :=$\vcentcolon=\{u\in H^{1}_{0}(\Omega)^{3}\,|\,u|_{K}\in P_{r+1}(K)^{3},\,\forall K\in\mathcal{T}_{h}\}$ (42a)
$U^{\text{div}}_{h}=RT_{r}(\mathcal{T}_{h})$ :=$\vcentcolon=\{B\in H_{0}(\operatorname{div},\Omega)\,|\,B|_{K}\in P_{r}(K)^{3}+xP_{r}(K),\,\forall K\in\mathcal{T}_{h}\}$ (42b)
$F_{h}=DG_{s}(\mathcal{T}_{h})$ :=$\vcentcolon=\{f\in L^{2}(\Omega)\,|\,f|_{K}\in P_{s}(K),\,\forall K\in\mathcal{T}_{h}\}.$ (42c)
@@
which satisfies $\operatorname{curl}U^{\operatorname{curl}}_{h}\subset U^{\text{div}}_{h}$. The $L^{2}(\Omega)$-orthogonal projections on $F_{h}$, $U^{\operatorname{div}}_{h}$, $U^{\operatorname{grad}}_{h}$, and $U^{\operatorname{curl}}_{h}$ are respectively denoted with $\pi_{h}$, $\pi^{\operatorname{div}}_{h}$, $\pi^{\operatorname{grad}}_{h}$, and $\pi^{\operatorname{curl}}_{h}$. Because discontinuous elements are used, some notations need to be introduced. Let $\mathcal{E}_{h}=\mathcal{E}^{0}_{h}\cup\mathcal{E}^{\delta}_{h}$ denote the set of $(d-1)$-dimensional faces in $\mathcal{T}_{h}$ with $\mathcal{E}^{0}_{h}$ the set of interior faces and $\mathcal{E}^{\delta}_{h}$ the set of boundary faces. For every face $e=K_{1}\cap K_{2}\in\mathcal{E}^{0}_{h}$, its length is written $h_{e}$, while the jump and average of a piecewise smooth scalar function $f$ are defined as
$\llbracket f\rrbracket$ $=f_{1}n_{1}+f_{2}n_{2}$ (44a)
@@
$H_{0}(\operatorname{curl},\Omega)$ =$=\{u\in L^{2}(\Omega)^{2}\,|\,\partial_{x}u_{y}-\partial_{y}u_{x}\in L^{2}(\Omega),\,u_{x}n_{y}-u_{y}n_{x}=0\text{ on }\partial\Omega\},$ (45a)
$U^{\text{curl}}_{h}$ =$=\{u\in H_{0}(\operatorname{curl},\Omega)\,|\,u|_{K}\in P_{r}(K)^{2}+(y,-x)P_{r}(K),\,\forall K\in\mathcal{T}_{h}\}.$ (45b)

### v2

@@ 1 Introduction
Dissipative magnetohydrodynamics. Plasma flows are at the core of numerous applications, including but not limited to: astrophysics, Magnetic Confinement Fusion [6, 28] and Inertial Confinement Fusion [2, 8]. Such flows display complex behaviors because of thermodynamics, interaction with the electromagnetic field, potentially highly compressible effects, and numerous and stiff irreversible processes. When characteristic velocities are negligible compared to the speed of light and the plasma is highly collisional, the flow can be reasonably described by Magnetohydrodynamics (MHD). Even for flows where the plasma is less collisional (i.e. for Magnetic Confinement Fusion), these equations are often used as a first approximation. The system under interest is
@@
with $n$ the outward normal vector to the boundary, together with a thermal boundary condition of either Dirichlet or Neumann type. Extension of these equations to include thermoelectric effect will be also considered. When dissipation (i.e. viscosity, resistivity and heat exchange) is removed, the so-called ideal MHD is recovered. It is well-known that ideal MHD equations can be derived from Hamilton’s principle [27]. Because dissipation is by nature irreversible and leads to an entropy production, it falls in the realm of non-equilibrium thermodynamics. An extension of Hamilton’s principle for non-equilibrium thermodynamics was developed in [15, 16]. It allows to derive equations automatically satisfying the two first laws of thermodynamics; namely conservation of energy and non-destruction of entropy. The above equations (1), including all irreversible processes, can be obtained from it, as will be shown shortly. This structure helps navigate through the otherwise complicated and intricate phenomena. More importantly, it will here serve as a guideline for deriving a numerical method that preserves key properties of the flow at the discrete level. Dissipative MHD has also been studied through extensions of Poisson brackets, including the metriplectic approach; see [26] and [7].
@@ 2.1 Ideal MHD
In the reversible case, as a consequence of mass and entropy conservation, and of the frozen-in property of magnetic induction, one has $\varrho(t)=\varrho_{0}$, $S(t)=S_{0}$ and $\mathcal{B}(t)=\mathcal{B}_{0}$ in the Lagrangian description. In this case, the MHD motion follows from Hamilton’s principle
@@
The 0-form $\gamma$ and $1$-form $Z$ will be used later in the irreversible case. The reduced Hamilton’s principle (4) then becomes
@@
$\partial_{t}\frac{\delta\ell}{\delta u}+\pounds_{u}\frac{\delta\ell}{\delta u}$ $=\rho\nabla\frac{\delta\ell}{\delta\rho}+s\nabla\frac{\delta\ell}{\delta s}+B\times\operatorname{curl}\frac{\delta\ell}{\delta B},$ (13a)
@@ 2.2 Visco-resistive MHD with heat conduction and thermoelectric effect
$\partial_{t}\frac{\delta\ell}{\delta u}+\pounds_{u}\frac{\delta\ell}{\delta u}=\rho\nabla\frac{\delta\ell}{\delta\rho}+s\nabla\frac{\delta\ell}{\delta s}+\operatorname{div}\sigma+B\times\operatorname{curl}\frac{\delta\ell}{\delta B},$ (22a)
$\overline{\mathfrak{D}}_{t}B=\operatorname{curl}j_{B},$ (22b)
$\overline{D}_{t}\rho=0,$ (22c)
$\displaystyle\frac{\delta\ell}{\delta s}(\overline{D}_{t}s+\operatorname{div}j_{s})=-\sigma:\nabla u-j_{s}\cdot\nabla\frac{\delta\ell}{\delta s}-j_{B}\cdot\operatorname{curl}\frac{\delta\ell}{\delta B},$ (22d)
@@
$j_{s}\cdot n=0,\quad j_{B}\times n=0,$ (23)
@@
The variational formulation given above yields thermally insulated boundaries, as we see from the first equation in (23), consistently with the variational setting of adiabatically closed systems, [15]. As shown in [13], a modification of the constraints (20a) and (20b), consistent with the variational principle for open systems, [17], allows the treatment of the Dirichlet boundary condition $T|_{\partial\Omega}=T_{0}$ or the nonhomogeneous Neumann boundary condition $j_{q}\cdot n=q_{0}$, with $j_{q}=Tj_{s}$ the heat flux, for given $T_{0},q_{0}:\partial\Omega\rightarrow\mathbb{R}$. To obtain these boundary conditions, the constraint (20a) is modified with the addition of a boundary term, namely,
$\int_{\Omega}w\frac{\delta\ell}{\delta s}\overline{D}_{t}\varsigma\,{\rm d}x=\int_{\Omega}w\big{(}-\sigma:\nabla u+j_{s}\cdot\nabla D_{t}\gamma+j_{B}\cdot\operatorname{curl}(\mathfrak{D}_{t}Z)\big{)}\,{\rm d}x-\int_{\partial\Omega}\!\!\!w(j_{s}\cdot n)\left(D_{t}\gamma-T_{0}\right)\,{\rm d}a,\;\forall\;w$ (27)
for the Dirichlet boundary conditions, and
$\int_{\Omega}w\frac{\delta\ell}{\delta s}\overline{D}_{t}\varsigma\,{\rm d}x=\int_{\Omega}w\big{(}-\sigma:\nabla u+j_{s}\cdot\nabla D_{t}\gamma+j_{B}\cdot\operatorname{curl}(\mathfrak{D}_{t}Z)\big{)}\,{\rm d}x-\int_{\partial\Omega}w\big{(}(j_{s}\cdot n)D_{t}\gamma-q_{0}\big{)}\,{\rm d}a,\;\forall\;w$ (28)
for the nonhomogeneous Neumann boundary condition. For both cases, the variational constraint (20b) is modified as
@@
The variational principle (19)–(21) with (20a) replaced by (27) or (28), and with (20b) replaced by (29), yield the same system (22)–(24), but with the boundary condition $j_{s}\cdot n=0$ replaced by $T|_{\partial\Omega}=T_{0}$ or $(j_{s}\cdot n)T=q_{0}$. In these cases, the energy balance in (25) is modified as
@@ 2.3 Physical closures
with $R$ the vector potential of the angular velocity, $\varepsilon(\rho,s)$ the internal energy density and $\mu_{0}>0$ the magnetic permeability. One then gets $\frac{\delta\ell}{\delta u}=\rho(u+R)$, $\frac{\delta\ell}{\delta\rho}=\frac{1}{2}|u|^{2}+R\cdot u-\frac{\partial\varepsilon}{\partial\rho}-\phi$, $\frac{\delta\ell}{\delta s}=\frac{\partial\varepsilon}{\partial s}=T>0$ the temperature, and $\frac{\delta\ell}{\delta B}=\frac{1}{\mu_{0}}B=H$. By defining the pressure $p=\frac{\partial\varepsilon}{\partial\rho}\rho+\frac{\partial\varepsilon}{\partial s}s-\varepsilon$, (22a) and (22d) eventually become
$\rho(\partial_{t}u+u\cdot\nabla u+2\omega\times u)$ $=-\rho\nabla\phi-\nabla p+\operatorname{curl}H\times B+\operatorname{div}\sigma,$ (32a)
@@
$j_{s}$ $=j_{s}(T)=-\frac{1}{T}\kappa\nabla T,$ $\text{with $\kappa\geq 0$},$ (33b)
$j_{B}$ $=j_{B}(H)=-\nu\operatorname{curl}H,$ $\text{with $\nu\geq 0$},$ (33c)
@@
$j_{s}$ $=j_{s}(T,H)=-\frac{1}{T}\kappa\nabla T+\alpha\operatorname{curl}H,$ $\text{with $\kappa\geq 0$},$ (35a)
$j_{B}$ $=j_{B}(T,H)=-\alpha\nabla T-\nu\operatorname{curl}H,$ $\text{with $\nu\geq 0$},$ (35b)
for some coefficient $\alpha$. Contrary to the previous coefficients, $\alpha$ is not subject to any sign restriction, as the thermoelectric effect only contributes to the entropy balance through a flux, the so-called Thomson effect. Note that if $\alpha$ is constant or a sole function of the temperature, it does not contribute to the induction equation.
@@ 2.4 Associated weak formulation
$e(w,f)=\left\{\begin{array}[]{ll}\vskip 5.69046pt\displaystyle 0&\text{for homogeneous Neumann}\\ \vskip-11.38092pt\displaystyle\int_{\partial\Omega}wj_{s}(f)\cdot nT_{0}\,{\rm d}a&\text{for Dirichlet}\\ \\ \displaystyle\int_{\partial\Omega}wq_{0}(f)\,{\rm d}a&\text{for nonhomogeneous Neumann}.\end{array}\right$. (38)
@@
With these definitions, the constraints (20a)–(20b), (27)–(29), and (28)–(29) can be written in a unified way for all boundary conditions as
@@
with $\left\langle\cdot,\cdot\right\rangle$ the $L^{2}$ inner product.
@@
By using these notations when carrying out the variational formulations (19)–(21) or (19)-(27)-(29)-(21) or (19)-(28)-(29)-(21), we get the following weak form of the equations (22) and associated boundary conditions $j_{s}\cdot n=0$ or $T|_{\partial\Omega}=T_{0}$ or $j_{q}\cdot n=q_{0}$ as follows
@@ 3.1 Discrete setting
$H^{1}_{0}(\Omega)$ $=\{f\in L^{2}(\Omega)\,|\,\nabla f\in L^{2}(\Omega)^{3},\,f=0\text{ on }\partial\Omega\},$ (42a)
$H_{0}(\operatorname{div},\Omega)$ $=\{u\in L^{2}(\Omega)^{3}\,|\operatorname{div}u\in L^{2}(\Omega),\,u\cdot n=0\text{ on }\partial\Omega\},$ (42b)
$H_{0}(\operatorname{curl},\Omega)$ $=\{u\in L^{2}(\Omega)^{3}\,|\,\operatorname{curl}u\in L^{2}(\Omega)^{3},\,u\times n=0\text{ on }\partial\Omega\}.$ (42c)
@@
$U^{\text{grad}}_{h}=CG_{r+1}(\mathcal{T}_{h})^{3}$ $\vcentcolon=\{u\in H^{1}_{0}(\Omega)^{3}\,|\,u|_{K}\in P_{r+1}(K)^{3},\,\forall K\in\mathcal{T}_{h}\}$ (43a)
$U^{\text{div}}_{h}=RT_{r}(\mathcal{T}_{h})$ $\vcentcolon=\{B\in H_{0}(\operatorname{div},\Omega)\,|\,B|_{K}\in P_{r}(K)^{3}+xP_{r}(K),\,\forall K\in\mathcal{T}_{h}\}$ (43b)
$F_{h}=DG_{s}(\mathcal{T}_{h})$ $\vcentcolon=\{f\in L^{2}(\Omega)\,|\,f|_{K}\in P_{s}(K),\,\forall K\in\mathcal{T}_{h}\}.$ (43c)
@@
which satisfies $\operatorname{curl}U^{\operatorname{curl}}_{h}\subset U^{\text{div}}_{h}$. The $L^{2}(\Omega)$-orthogonal projections on $F_{h}$, $U^{\operatorname{div}}_{h}$, $U^{\operatorname{grad}}_{h}$, and $U^{\operatorname{curl}}_{h}$ are respectively denoted with $\pi_{h}$, $\pi^{\operatorname{div}}_{h}$, $\pi^{\operatorname{grad}}_{h}$, and $\pi^{\operatorname{curl}}_{h}$. Because discontinuous elements are used, some notations need to be introduced. Let $\mathcal{E}_{h}=\mathcal{E}^{0}_{h}\cup\mathcal{E}^{\delta}_{h}$ denote the set of codimension-$1$ faces in $\mathcal{T}_{h}$ with $\mathcal{E}^{0}_{h}$ the set of interior faces and $\mathcal{E}^{\delta}_{h}$ the set of boundary faces. For every face $e=K_{1}\cap K_{2}\in\mathcal{E}^{0}_{h}$, its length is written $h_{e}$, while the jump and average of a piecewise smooth scalar function $f$ are defined as
$\llbracket f\rrbracket$ $=f_{1}n_{1}+f_{2}n_{2},$ (45a)
@@
$H_{0}(\operatorname{curl},\Omega)$ $=\{u\in L^{2}(\Omega)^{2}\,|\,\partial_{x}u_{y}-\partial_{y}u_{x}\in L^{2}(\Omega),\,u_{x}n_{y}-u_{y}n_{x}=0\text{ on }\partial\Omega\},$ (46a)
$U^{\text{curl}}_{h}$ $=\{u\in H_{0}(\operatorname{curl},\Omega)\,|\,u|_{K}\in P_{r}(K)^{2}+(y,-x)P_{r}(K),\,\forall K\in\mathcal{T}_{h}\}.$ (46b)